OpenStudy (amberlykhan):

Are there any values of x besides 2 in which x^2 = 2x?

OpenStudy (dariusx):

0

OpenStudy (dariusx):

x^2 - 2x = 0 x(x-2)=0 x = 0 and x -2 =0 x = 2